OpenStudy (anonymous):

Sketch the graph of the function below, being sure to include at least two full periods. y=sin(x+pi/4). Here's what I am starting to understand? So, my amp =3/4 and period of one cycle is 2pi/1?

sam (.sam.):

amplitude is 1 , y=1sin(x+pi/4)

OpenStudy (anonymous):

@sam...I kinda wrote that wrong..

OpenStudy (anonymous):

I looked at the other problem above the one I shared. I wanted to share y=3/4 sin (2pi-pi).

OpenStudy (anonymous):

My amplitude is 2pi/1?

sam (.sam.):

\[y=\frac{3}{4} \sin (2\pi-\pi)\] This one?

OpenStudy (anonymous):

YES! Thanks. That's it!

OpenStudy (anonymous):

:)

sam (.sam.):

are you sure its sin(2PI-PI)?

OpenStudy (anonymous):

lol. It is sin(2x-pi)....thanks for hanging in there.

sam (.sam.):

\[y=\frac{3}{4} \sin (2x-\pi)\] So the amplitude is 3/4

OpenStudy (anonymous):

I agree. My period is? That's the part that is hard to see just by looking at it. Just wondering if there's a way to do that because I need to learn to do this more quickly.

OpenStudy (anonymous):

would I look at the -pi for the amplitude? I have seen this done two different ways. I am trying to make it work for me.

sam (.sam.):

Formula \[y = Asin(bx + c) + d\] \[T=\frac{2\pi}{b}\]

sam (.sam.):

So, the period ,\[T=\frac{2\pi}{2}=\pi\]

OpenStudy (anonymous):

how did you know to use those numbers.

sam (.sam.):

The formula to find period is \(T=\frac{2\pi}{b}\) ------------------------------------------- So, \[y=\frac{3}{4} \sin (2x-\pi) \\ \\ \Big\downarrow \\ \\ y=Asin(bx+c)\] --------------------------------------------- b is 2, \[T=\frac{2\pi}{2}=\pi\]

OpenStudy (anonymous):

cool...thanks!

OpenStudy (anonymous):

Ok, so here's an example of how my prof. does this.

OpenStudy (anonymous):

y=sin(x+pi/4) (no period, no change no amp)..maybe a bad example. The best I could find .

OpenStudy (anonymous):



OpenStudy (anonymous):